if i am given the radius of a circle which is 3 and the cylinder has a lateral SA of 60 what is the height?... (i have tried many diff. things and cant get it right)
Answer by stanbon(75887)

if i am given the radius of a circle which is 3 and the cylinder has a lateral SA of 60 what is the height?...
---
The lateral SA = (circumference*height)
----
SA = (2(pi)r)(height)
----
60 = (2(pi)*3)h
---
h = 10/pi
----
height ~ 3.183 units
=====================================
Cheers,
Stan H.